OpenStudy (anonymous):

(4x-1)(-2x-3)+(1-4x)(3x+2)+16x^2=1

OpenStudy (lncognlto):

Can I show you how I would solve the question?

OpenStudy (anonymous):

As long it is right, do as you wisch

OpenStudy (anonymous):

*wish.

OpenStudy (lncognlto):

Right then, using the FOIL method on (4x-1)(-2x-3) gives us: -8x^2 - 12x + 2x +3 =-8x^2 -10x + 3.

OpenStudy (anonymous):

I have got the same.

OpenStudy (lncognlto):

Good, now using it on (1-4x)(3x+2) gives: 3x + 2 - 12x^2 - 8x. = - 12x^2 - 5x + 2.

OpenStudy (anonymous):

5(-4x^2-3x+1)

OpenStudy (lncognlto):

Now the equation looks like: -8x^2 -10x + 3 - 12x^2 - 5x + 2 + 16x^2 = 1 therefore -4x^2 - 15x + 5 = 1 tf 4x^2 + 15x - 4 = 0

OpenStudy (lncognlto):

tf 4x^2 + 16x - x - 4 = 0 tf 4x(x + 4) - 1(x + 4) = 0 tf (4x -1)(x+4) = 0 tf 4x -1 = 0 or x + 4 = 0 tf x = 1/4 or x = -4

OpenStudy (anonymous):

Thank you.

OpenStudy (lncognlto):

You're welcome :)